# Understanding Flow Behaviors of Supercooled Liquids by Embodying Solid-Liquid Duality at Particle Level

### v1

@@ V.2 Steady shear
Here, the subscript “sd” denotes “shear-driven”, meaning that the local configuration and relaxation of particles corresponding to the gaussian-like component are strongly driven by shear. We plot the results of $p_{\mathrm{sd}}$ as a function of $\dot{\gamma}$ for two samples in Fig. 27 (c) and (d). According to the above analysis, the gaussian-like component indicates slow particles. Thus, in Fig. 27 (c) and (d), we replot the results of $p_{\mathrm{slow}}$ shown in Fig. 13, which is obtained from the MD data, for comparison. Their dependences on $\dot{\gamma}$ are seen to be similar. Especially, $p_{\mathrm{sd}}$ gradually saturates as the system enters the shear-thinning regime. The observations found in Fig. 27 suggest that the SGR model exhibits two groups of responses to shear, which captures the major feature of the grouping of slow/fast particles.
Figure 28: Normalized shear viscosity $\eta^{*}$ as a function of $\dot{\gamma}$. (a) and (b) show the results of sample A and sample B, respectively. Dashed lines denote the SGR results calculated with Eq. 32. Symbols denote the results found from MD data.
By adopting the strategy that the macroscopic stress is proportional to the average of the atomic level stress [69], we can evaluate the shear stress of the system by the stress experienced by a particle in its basin. For simplicity, this atomic level stress can be calculated by $\mathrm{d}E(\gamma)/\mathrm{d}\gamma$, where $E(\gamma)$ is given by Eq. 16. Therefore, under steady state, the shear stress derived from the SGR model is given by:
@@
Then, the SGR viscosity can be found by $\eta_{\mathrm{SGR}}=\sigma_{\mathrm{SGR}}/\dot{\gamma}$. In Fig. 28, we show the results of $\eta_{\mathrm{SGR}}$, normalized as $\eta_{\mathrm{SGR}}^{*}=\eta_{\mathrm{SGR}}/\eta_{\mathrm{eq}}$, for two samples. $\eta_{\mathrm{SGR}}(\dot{\gamma})$ captures the general features of shear thinning: As $\dot{\gamma}$ increases, it is able to delineate three rheological regimes, i.e., Newtonian regime, crossover regime, and shear-thinning regime. Compared with Fig. 27 (c) and (d), it can be seen that the crossover regime appears when $p_{\mathrm{sd}}$ becomes finite, and the shear-thinning regime appears when $p_{\mathrm{sd}}$ starts to saturate. This result suggests that the emergence of shear thinning is due to the increasingly dominant role of shear-driven particles in the system. The picture given by the PEL framework is consistent with our MD results shown in Sec. III.
We also plot the MD results of the normalized viscosity in Fig. 28. The SGR results and the MD results exhibit similar behaviors at the qualitatively level. Nevertheless, the quantitative deviation appears and grows as $\dot{\gamma}$ increases. We will discuss about this point in the next subsection.
To briefly summarize these two subsections, we try to rationalize the MD results found in Secs. IIIV, which connect microscopic structure to macroscopic flow behaviors of supercooled liquids, within the framework of PEL. The PEL approach simplifies the intricate many-body interaction into particle-level basins. Moreover, it condenses the structural information characterized by $g_{0,i}^{0}$ and $g_{2,i}^{-2}$ into the height and strain characteristics of basins, enabling us to quantify the connection between local configuration and dynamics. The physical significance of $\tau_{\mathrm{LC}}$ and $\tau_{\mathrm{sh}}$, and the effects of shear flow and thermal activation, are clearly illustrated in the PEL picture. With the aid of the SGR model, the relation between structure and rheology in flow state is clarified. Particularly, the model exhibits two distinct shear responses as evidenced by the bimodal distribution shown in Fig. 27 (a) and (b). The relative proportion of these two responses dictates the variation of viscosity with respect to $\dot{\gamma}$, as illustrated in Fig. 28.
@@ V.3 Correlation
The systematic deviation between model results and MD results shown in Fig. 28 could be due to many reasons. For example, the basin shape $E(\gamma)$ we adopt could be oversimplified [142]; the restructuring process is ignored in the SGR model, etc. In our opinion, one of the key reasons is that the model given by Eqs. 1921 ignores the correlation among the cage-jump events of different particles. In fact, it has been demonstrated that there exists an elasticity-mediated long-ranged correlation among local flow events in glasses [92, 108]. In our recent work [83], we also find the evidence of dynamic facilitation mediated by elasticity in solid-like regions in flowing supercooled liquids. In principle, as $\dot{\gamma}$ increases, such long-ranged correlation in a supercooled liquid should become more significant as the solid-like component grows. To account for this correlation while maintaining the simplicity of the model, one can replace the thermal term $k_{\mathrm{B}}T$ in Eq. 19 by the effective activation temperature $x$ [23]. Although the use of an effective temperature to describe the effect of mechanical noise has been questioned [143], it can qualitatively capture the phenomenon that the correlation among particles can facilitate the relaxation of a region.
@@
Figure 29: (a) $\bar{\tau}_{\mathrm{sh}}(\dot{\gamma})$ (green circles, found from MD data) and $\bar{t}(\dot{\gamma})$ (defined in Eq. 33) for sample B. Here, $\bar{t}(\dot{\gamma})$ are calculated in two ways: with constant $\Omega$ (blue points) and with varying $\Omega$ determined by Eq. 34 (orange points). Inset: the varying $\Omega$ (red circles), determined by Eq. 34, as a function of $\dot{\gamma}$. The dashed line denotes the linear fit of $\Omega$ with respect to $\dot{\gamma}$. (b) Normalized viscosity of sample B. Symbols denote the MD result. Black dashed line denotes the result calculated with Eq. 32 with constant $\Omega$. This result has been shown in Fig. 28 (b). Orange dashed line denotes the result calculated with the varying $\Omega$ shown in the inset of (a).
We use the following strategy to determine $\Omega$. In Sec. V.1, we introduce $\bar{t}(E_{\mathrm{b}},\dot{\gamma})$ in Eq. 18 to represent $\tau_{\mathrm{sh}}$ in the start-up shear. The results are shown in Fig. 23. It is seen that the values of $\tau_{\mathrm{sh}}$ evaluated by $\bar{t}(E_{\mathrm{b}},\dot{\gamma})$ and those measured from MD data have small but non-negligible discrepancy. Thus, we can determine a reasonable $\Omega$ by minimizing the difference between $\bar{t}(E_{\mathrm{b}},\dot{\gamma})$ and $\tau_{\mathrm{sh}}$ for each $\dot{\gamma}$. To be specific, we introduce the basin-averaged $\bar{t}(E_{\mathrm{b}},\dot{\gamma})$, denoted as $\bar{t}(\dot{\gamma})$:
@@
In Fig. 29 (a), we present $\bar{\tau}_{\mathrm{sh}}(\dot{\gamma})$ and $\bar{t}(\dot{\gamma})$ using sample B as an illustrative example. Here, $\bar{t}(\dot{\gamma})$ is computed with two ways: the blue symbols give the result calculated with a constant $\Omega$; the orange symbols give the result calculated with the $\dot{\gamma}$-dependent $\Omega$ determined by Eq. 34. Generally speaking, incorporating a constant $\Omega$ leads to an over-estimation of $\tau_{\mathrm{sh}}$ at large $\dot{\gamma}$. In the inset of Fig. 29 (a), we give the result of $\Omega(\dot{\gamma})$ determined by Eq. 34 for sample B. $\Omega(\dot{\gamma})$ displays an increasing trend with $\dot{\gamma}$, corresponding to the enhancement of the correlation among flow events due to the increasingly dominant role of solid-like regions as $\dot{\gamma}$ increases. Within the mean-field framework of SGR, the increment of $\Omega(\dot{\gamma})$ can be understood by the enhancement of the rigidity [144]. As the system becomes more rigid, the relaxation of the central particle will influence more surrounding particles, resulting in a larger $\Omega$ [15, 80].
By replacing the constant $\Omega$ in Eq. 32 with $\Omega(\dot{\gamma})$, we recalculate the viscosity, and plot the result in Fig. 29 (b). We also replot the MD result and the result calculated by Eq. 32 with the constant $\Omega$, which have been shown in Fig. 28 (b). It is seen that by using $\Omega(\dot{\gamma})$, the agreement between MD result and model result is significantly improved.
@@ VI CONCLUDING REMARKS
The key to bridging the gap between the microscopic and the macroscopic worlds is the local configurational relaxation time $\tau_{\mathrm{LC}}$. Defined at the particle level, $\tau_{\mathrm{LC}}$ well reflects the local mobility and the heterogeneous feature of dynamics. Moreover, $\tau_{\mathrm{LC}}$ plays a role similar to that of $\tau_{\mathrm{M}}$ in the Maxwell model: Depending on the comparison between $\tau_{\mathrm{LC}}$ and the external driving rate, the corresponding local region can exhibit solid-like or liquid-like response. Then, the macroscopic behavior is given by the average behavior of all local regions in the system. These results unify the concept of solid-liquid duality and DH into a micro-mechanical picture for the flow of supercooled liquids. $\tau_{\mathrm{LC}}$ has an unambiguous structural foundation. With the aid of statistical learning methods and PEL, we show that $\tau_{\mathrm{LC}}$ is mainly determined by the combined effect of local packing and local configurational distortion, which are respectively represented by the barrier height and the tilted strain of the particle-level energy basin. With these efforts, we establish a chain relation of local configuration – local dynamics – local response – flow behaviors for supercooled liquids.
@@ Acknowledgements.
This research was partially supported by the National Natural Science Foundation of China (no. 11975136). The authors acknowledge the Center of High Performance Computing, Tsinghua University for providing computational resources and Dr. Dejia Kong for his help in Sec. IV.
Appendix A CIRCLING CLUSTERS
@@
Figure 30: Illustration of the “Grids” method for identification of slow/fast clusters. Here, the system is divided into grids. Each grid is assigned a value, which can be $\tau_{\mathrm{LC}}$ or other variables. Without loss of generality, the values of the grids are normalized to the range of 00 to $1$. Giving a typical example, we define the regions of interest as those grids whose values are greater than $0.5$. The boundaries of these regions can be identified. Here, we use points to represent the boundaries. Our objective is to divide the regions of interest into clusters. First, we identify the center grids. A center grid is defined as the grid whose value is greater than those of all its neighboring grids. In this figure, cross symbols are used to denote the centers. Subsequently, each grid within the region of interest is assigned to the nearest center grid. As a result, each center grid has its “jurisdiction area”, which is represented by solid lines in the figure.
As seen from Fig. 2, the clusters of slow particles exhibit hierarchical characteristics. Thus, it can be imagined that slow particles “grow” from different slow centers and form different slow clusters. Consequently, we can partition the particles based on the spatial positions of the slow centers. The identification of slow centers is carried out using the “Grids” method. In brief, the simulation box is divided into small cubic grids. The side length of each grid is $L/20$, where $L$ is the length of the simulation box. Subsequently, with the following expression, each grid undergoes a convolution with particles’ $\tau_{\mathrm{LC}}$, yielding a value denoted as $g$:
@@
where the subscript “$i$” denotes the $i$th grid, $\tau_{\mathrm{LC},j}$ is the $\tau_{\mathrm{LC}}$ of particle $j$, $r_{ij}$ is the distance between particle $j$ and grid $i$, and $l_{\mathrm{cg}}$ is a characteristic length employed in the convolution and is set to $L/20$. The summations in Eq. 35 are performed over all particles.
We define slow centers as the grids whose $g$ values are greater than those of their $26$ neighboring grids. This criterion guarantees that the slow centers are identified as local maxima. By utilizing these slow centers, each slow particle is assigned to the nearest center, thereby enabling the division of all slow particles into distinct clusters. The schematic diagram of this methods is presented in Fig. 30.

### v2

@@ V.2 Steady shear
Here, the subscript “sd” denotes “shear-driven”, meaning that the local configuration and relaxation of particles corresponding to the gaussian-like component are strongly driven by shear. We plot the results of $p_{\mathrm{sd}}$ as a function of $\dot{\gamma}$ for two samples in figure 27 (c) and (d). In figure 27 (c) and (d), we replot the results of $p_{\mathrm{slow}}$ shown in figure 13, which is obtained from the MD data, for comparison. Their dependences on $\dot{\gamma}$ are seen to be similar. Especially, $p_{\mathrm{sd}}$ gradually saturates as the system enters the shear-thinning regime. Figure 27 suggests that the SGR model exhibits two groups of responses to shear, which captures the major feature of the grouping of slow/fast particles.
Figure 28: Normalized shear viscosity $\eta^{*}$ as a function of $\dot{\gamma}$. (a) and (b) show the results of sample A and sample B, respectively. Dashed lines denote the SGR results calculated with equation 32. Symbols denote the results found from MD data.
Knowing that the macroscopic stress is given by the average of the atomic level stress [69], we evaluate the shear stress of the system with the stress experienced by a particle in its basin. For simplicity, this atomic level stress can be calculated by $\mathrm{d}E(\gamma)/\mathrm{d}\gamma$, where $E(\gamma)$ is given by equation 16. Therefore, under steady state, the shear stress derived from the SGR model is given by:
@@
Then, the SGR viscosity can be found by $\eta_{\mathrm{SGR}}=\sigma_{\mathrm{SGR}}/\dot{\gamma}$. In figure 28, we show the results of $\eta_{\mathrm{SGR}}$, normalized as $\eta_{\mathrm{SGR}}^{*}=\eta_{\mathrm{SGR}}/\eta_{\mathrm{eq}}$, for two samples. $\eta_{\mathrm{SGR}}(\dot{\gamma})$ captures the general features of shear thinning: As $\dot{\gamma}$ increases, it is able to delineate three rheological regimes, i.e., Newtonian regime, crossover regime, and shear-thinning regime. Compared with figure 27 (c) and (d), it can be seen that the crossover regime appears when $p_{\mathrm{sd}}$ becomes finite, and the shear-thinning regime appears when $p_{\mathrm{sd}}$ starts to saturate. This result suggests that the emergence of shear thinning is due to the increasingly dominant role of shear-driven particles in the system. The picture given by the model analysis is consistent with our MD results shown in section III.
We also plot the MD results of the normalized viscosity in figure 28. The SGR results and the MD results exhibit similar behaviors at the qualitatively level. Nevertheless, the quantitative deviation appears and grows as $\dot{\gamma}$ increases. We will discuss about this point in the next subsection.
To briefly summarize these two subsections, we try to model the MD results found in sections IIIV, which connect microscopic structure to macroscopic flow behaviors of supercooled liquids, by considering the shear-facilitated activation from basins. This picture condenses the structural information characterized by $g_{0,i}^{0}$ and $g_{2,i}^{-2}$ into the height and strain characteristics of particle-level basins, enabling us to quantify the connection between local configuration and dynamics. The physical significance of $\tau_{\mathrm{LC}}$ and $\tau_{\mathrm{sh}}$, and the effects of shear flow and thermal activation, are clearly illustrated in this picture. With the aid of the SGR model, the relation between structure and rheology in flow state is clarified. Particularly, the model exhibits two distinct shear responses as evidenced in figure 27. The relative proportion of these two responses dictates the variation of viscosity with respect to $\dot{\gamma}$, as illustrated in figure 28.
@@ V.3 Correlation
The deviation between model results and MD results shown in figure 28 could be due to many reasons. For example, the basin shape $E(\gamma)$ we adopt could be oversimplified [141]; the restructuring process is ignored in the SGR model, etc. In our opinion, one of the key reasons is that the model given by equations 1921 ignores the correlation among the cage-jump events of different particles. It has been demonstrated that there exists an elasticity-mediated long-ranged correlation among local flow events in glasses [92, 108]. In our recent work [83], we also find the evidence of dynamic facilitation mediated by elasticity in solid-like regions in flowing supercooled liquids. In principle, as $\dot{\gamma}$ increases, such long-ranged correlation in a supercooled liquid should become more significant as the solid-like component grows. To account for this correlation while maintaining the simplicity of the model, one can replace the thermal term $k_{\mathrm{B}}T$ in equation 19 by the effective activation temperature $x$ [23]. Although the use of an effective temperature to describe the effect of mechanical noise has been questioned [133], it can qualitatively capture the phenomenon that the correlation among particles can facilitate the relaxation of a region.
@@
Figure 29: (a) $\bar{\tau}_{\mathrm{sh}}(\dot{\gamma})$ (green circles, found from MD data) and $\bar{t}(\dot{\gamma})$ (defined in equation 33) for sample B. Here, $\bar{t}(\dot{\gamma})$ are calculated in two ways: with constant $\Omega$ (blue points) and with varying $\Omega$ determined by equation 34 (orange points). Inset: the varying $\Omega$ (red circles), determined by equation 34, as a function of $\dot{\gamma}$. The dashed line denotes the linear fit of $\Omega$ with respect to $\dot{\gamma}$. (b) Normalized viscosity of sample B. Symbols denote the MD result. Black dashed line denotes the result calculated with equation 32 with constant $\Omega$. This result has been shown in figure 28 (b). Orange dashed line denotes the result calculated with the varying $\Omega$ shown in the inset of (a).
We use the following strategy to determine $\Omega$. In section V.1, we introduce $\bar{t}(E_{\mathrm{b}},\dot{\gamma})$ in equation 18 to represent $\tau_{\mathrm{sh}}$ in the start-up shear. The results are shown in figure 23. It is seen that the values of $\tau_{\mathrm{sh}}$ evaluated by $\bar{t}(E_{\mathrm{b}},\dot{\gamma})$ and those measured from MD data have small but non-negligible discrepancy. Thus, we can determine a reasonable $\Omega$ by minimizing the difference between $\bar{t}(E_{\mathrm{b}},\dot{\gamma})$ and $\tau_{\mathrm{sh}}$ for each $\dot{\gamma}$. To be specific, we introduce the basin-averaged $\bar{t}(E_{\mathrm{b}},\dot{\gamma})$, denoted as $\bar{t}(\dot{\gamma})$:
@@
In figure 29 (a), we present $\bar{\tau}_{\mathrm{sh}}(\dot{\gamma})$ and $\bar{t}(\dot{\gamma})$ using sample B as an illustrative example. Here, $\bar{t}(\dot{\gamma})$ is computed with two ways: the blue symbols give the result calculated with a constant $\Omega$; the orange symbols give the result calculated with the $\dot{\gamma}$-dependent $\Omega$ determined by equation 34. Generally speaking, incorporating a constant $\Omega$ leads to an over-estimation of $\tau_{\mathrm{sh}}$ at large $\dot{\gamma}$. In the inset of figure 29 (a), we give the result of $\Omega(\dot{\gamma})$ determined by equation 34 for sample B. $\Omega(\dot{\gamma})$ displays an increasing trend with $\dot{\gamma}$, corresponding to the enhancement of the correlation among flow events due to the increasingly dominant role of solid-like regions as $\dot{\gamma}$ increases. Within the mean-field framework of SGR, the increment of $\Omega(\dot{\gamma})$ can be understood by the enhancement of the rigidity [142]. As the system becomes more rigid, the relaxation of the central particle will influence more surrounding particles, resulting in a larger $\Omega$ [15, 80].
By replacing the constant $\Omega$ in equation 32 with $\Omega(\dot{\gamma})$, we recalculate the viscosity, and plot the result in figure 29 (b). We also replot the MD result and the result calculated by equation 32 with the constant $\Omega$, which have been shown in figure 28 (b). It is seen that by using $\Omega(\dot{\gamma})$, the agreement between MD result and model result is significantly improved.
@@ VI CONCLUDING REMARKS
The key to bridging the gap between the microscopic and the macroscopic worlds is the local configurational relaxation time $\tau_{\mathrm{LC}}$. Defined at the particle level, $\tau_{\mathrm{LC}}$ well reflects the local mobility and the heterogeneous feature of dynamics. Moreover, $\tau_{\mathrm{LC}}$ plays a role similar to that of $\tau_{\mathrm{M}}$ in the Maxwell model: Depending on the comparison between $\tau_{\mathrm{LC}}$ and the external driving rate, the corresponding local region can exhibit solid-like or liquid-like response. Then, the macroscopic behavior is given by the average behavior of all local regions in the system. These results unify the concept of solid-liquid duality and DH into a micro-mechanical picture for the flow of supercooled liquids. $\tau_{\mathrm{LC}}$ has an unambiguous structural foundation. With the aid of statistical learning methods and a model considering the shear-facilitated activation from energy basins, we show that $\tau_{\mathrm{LC}}$ is determined by the combined effect of local packing and local configurational distortion, which are respectively represented by the barrier height and the tilted strain of the particle-level energy basin. With these efforts, we establish a chain relation of local configuration – local dynamics – local response – flow behaviors for supercooled liquids.
As for the solid-liquid duality, extensive studies of amorphous matter, including but not limited to the glass transition and flow of supercooled liquids and the yielding of amorphous solids, are built on this intuition, with concepts such as soft modes [27, 84], shear transformation zones [86], cooperative rearranging regions [45] and softness [30]. Most of these studies focus on the soft, liquid-like regions. On the other hand, our study highlights the increasing importance of the hard, solid-like regions in determining the flow behaviors of supercooled liquids as the flow is enhanced. This difference represents a shift of consideration [44]: For the yielding of solids, the key question is why a solid fails or even flows like a liquid. Liquid-like regions (for example, STZ) in the solid background naturally play as the precursor of the bulk yielding. For the flow of supercooled liquids, on the contrary, the key question is shifted to why a liquid exhibits strong viscoelasticity. By highlighting the solid-like regions in the liquid background, we are able to quantify the significant nonlinear behaviors of the flow.
@@ Acknowledgements.
This research was partially supported by the National Natural Science Foundation of China (no. 11975136). The authors acknowledge the Center of High Performance Computing, Tsinghua University for providing computational resources and Dr. Dejia Kong for his help in section IV.
Appendix A ALGORITHMS FOR IDENTIFYING CAGE JUMPS AND CIRCLING CLUSTERS
A.1 Cage jump
After obtaining the trajectory $S(t)$ of a particle from 0 to $t_{\text{end}}$, (in shear condition, trajectory used is the non-affine trajectory, where the pure shear influence has been eliminated [45]), an iterative process can be carried out to find cage jumps [59, 60]. Specifically, the trajectory is divided at an intermediate time $t_{\mathrm{m}}$, resulting in two sub-trajectories: $d_{1}(t)$ from 0 to $t_{\mathrm{m}}$, and $d_{2}(t)$ from $t_{\mathrm{m}}$ to $t_{\text{end}}$. To evaluate the degree of separation between $d_{1}$ and $d_{2}$ at $t_{\mathrm{m}}$, we calculate
$p(t_{\mathrm{m}})=\zeta(t_{\mathrm{m}})\sqrt{\big{\langle}d_{1}(t)-\overline{d}_{2}(t)\big{\rangle}^{2}_{d_{1}}\cdot\big{\langle}d_{2}(t)-\overline{d}_{1}(t)\big{\rangle}^{2}_{d_{2}}},$ (35)
where $\zeta(t_{\mathrm{m}})=\sqrt{t_{\mathrm{m}}/t_{\text{end}}\times(1-t_{\mathrm{m}}/t_{\text{end}})}$, $\overline{d}(t)$ represents the center of mass of the trajectory $d(t)$, and $\langle\cdot\cdot\cdot\rangle_{d(t)}$ denotes the average over the trajectory $d(t)$. This evaluation is performed for every time point along the trajectory to monitor when the $p(t_{\mathrm{m}})$ reaches its maximum value $p_{\text{max}}$, and the corresponding time $t_{\mathrm{m,max}}$. Subsequently, the same process is applied to each sub-trajectory—from 0 to $t_{\mathrm{m,max}}$ and $t_{\mathrm{m,max}}$ to $t_{\text{end}}$—to determine the critical times for these sub-trajectories. This division can be recursively applied to further sub-trajectories until $p_{\text{max}}$ falls below a threshold that, represents the cage size. Through this iterative procedure, cage jumps in the trajectory can be identified when $p(t)$ is maximal.
A.2 Circling clusters
@@
As seen from figure 2, the clusters of slow particles exhibit hierarchical characteristics. Thus, it can be imagined that slow particles “grow” from different slow centers and form different slow clusters. Consequently, we can partition the particles based on the spatial positions of the slow centers. The identification of slow centers is carried out using the “Grids” method. In brief, the simulation box is divided into small cubic grids. The side length of each grid is $L/20$, where $L$ is the length of the simulation box. Subsequently, with the following expression, each grid undergoes a convolution with particles’ $\tau_{\mathrm{LC}}$, yielding a value denoted as $g$:
@@
where the subscript “ii” denotes the $i$th grid, $\tau_{\mathrm{LC},j}$ is the $\tau_{\mathrm{LC}}$ of particle $j$, $r_{ij}$ is the distance between particle $j$ and grid $i$, and $l_{\mathrm{cg}}$ is a characteristic length employed in the convolution and is set to $L/20$. The summations in equation 36 are performed over all particles.
We define slow centers as the grids whose $g$ values are greater than those of their $26$ neighboring grids. This criterion guarantees that the slow centers are identified as local maxima. By utilizing these slow centers, each slow particle is assigned to the nearest center, thereby enabling the division of all slow particles into distinct clusters. The schematic diagram of this methods is presented in figure 30.
Figure 30: Illustration of the “Grids” method for identification of slow/fast clusters. Here, the system is divided into grids. Each grid is assigned a value, which can be $\tau_{\mathrm{LC}}$ or other variables. Without loss of generality, the values of the grids are normalized to the range of 0 to $1$. Giving a typical example, we define the regions of interest as those grids whose values are greater than $0.5$. The boundaries of these regions can be identified. Here, we use points to represent the boundaries. Our objective is to divide the regions of interest into clusters. First, we identify the center grids. A center grid is defined as the grid whose value is greater than those of all its neighboring grids. In this figure, cross symbols are used to denote the centers. Subsequently, each grid within the region of interest is assigned to the nearest center grid. As a result, each center grid has its “jurisdiction area”, which is represented by solid lines in the figure.
Figure 31: (a) and (b) show the relation between the average neighbor-keeping ratio among slow particles, denoted as $\bar{C}$, and the ratio of particles that have experienced cage jump in the system, denoted as $\phi$, for sample A and sample B, respectively, at various $\dot{\gamma}$. (c) and (d) show the results of $w$ (symbols), found from equation 39, as a function of $\dot{\gamma}$ for sample A and sample B, respectively. In (c) and (d), horizontal dashed lines denote the mean values of $w$ averaged over all $\dot{\gamma}$ for sample A and sample B.